# Gauge Flow Models

### v2

@@ A Gauge Flow Model is defined within the mathematical framework of associated bundles:
  • $\Pi_{M}:P\times F\to TM$ is a smooth, learnable projection map.
@@ 2.1  Fiber Bundle
A fiber bundle is a geometric construction that assembles a continuous family of fibers, one for each point of a base space, in a way that varies continuously across the base. Locally, it resembles a product of the base space and the fiber. However, it’s important to note that fiber bundles can generally deviate from this local product space in non-trivial ways. An illustration of a fiber bundle is provided below.
@@
$\pi^{-1}(U_{x})$$U\times F$$U_{x}$$U_{x}$$\scriptstyle{\Psi}$$\scriptstyle{\pi}$$\scriptstyle{\mathrm{pr}_{1}}$$\scriptstyle{\mathrm{id}_{U_{x}}}$
@@ 2.1.2  Principal Bundle
    $g\bullet g^{-1}=e=g^{-1}\bullet g\quad\exists g^{-1}\in G,\;\ \forall g\in G$
@@
where the section is defined such that its projection onto $U\subset M$ recovers the identity on $U$:
@@ A Gauge Flow Model is defined within the framework of the associated bundle $\hat{A}$:
  • $\Pi_{M}:P\times F\rightarrow TM$ is a smooth projection from the product bundle $P\times F$ to the tangent bundle $TM$ of the base manifold $M$. This projection is covered, ensuring that the projection of any point $(p,v)$ from the product bundle collapses to the same base point $p$ within $M$:
    $\pi_{TM}(\Pi_{M}(p,v))=\pi_{P}(p)\quad\forall(p,v)\in P\times F$
@@
$\mathcal{L}_{\mathrm{GFM}}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ x\sim p_{t}\end{subarray}}\bigg{\lVert}\Big{[}v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu}(x(t),t)d^{\mu}(x(t),t)v_{s}(x(t),t))\Big{]}-u_{t}(x)\bigg{\rVert}_{g_{x}}^{2}$
@@ 4.1  Model
  • The mapping $\Pi_{M}:P\times\mathbb{R}^{N}\rightarrow M\times\mathbb{R}^{N}$ acts as the identity function on both $M$ and $\mathbb{R}^{N}$, with the $G$-component effectively projected out as a consequence of the chosen gauge.

### v3

@@ A Gauge Flow Model is defined within the mathematical framework of associated bundles:
  • $\Pi_{M}:P\times_{G}F\to TM$ is a smooth, learnable projection map.
@@ 2.1  Fiber Bundle
A fiber bundle is a geometric construction that assembles a continuous family of fibers, one for each point of a base space, in a way that varies continuously across the base. Locally, it resembles a product of the base space and the fiber. However, it’s important to note that while they are locally trivial, fiber bundles can generally deviate from being a global product space in non-trivial ways. An illustration of a fiber bundle is provided below.
@@
$\pi^{-1}(U_{x})$$U_{x}\times F$$U_{x}$$U_{x}$$\scriptstyle{\Psi}$$\scriptstyle{\pi}$$\scriptstyle{\mathrm{pr}_{1}}$$\scriptstyle{\mathrm{id}_{U_{x}}}$
@@ 2.1.2  Principal Bundle
    $g\bullet g^{-1}=e=g^{-1}\bullet g\quad\forall g\in G,\;\ \exists g^{-1}\in G$
@@
where the section is defined such that its projection onto $U\subset B$ recovers the identity on $U$:
@@ A Gauge Flow Model is defined within the framework of the associated bundle $\hat{A}$:
  • $\Pi_{M}:P\times_{G}F\rightarrow TM$ is a smooth projection map from the associated bundle $\hat{A}=P\times_{G}F$ to the tangent bundle $TM$. This map covers the identity on $M$, ensuring that any element $p,v]\in P\times_{G}F$ is projected to the tangent space of its corresponding base point in $M$:
    $\pi_{TM}(\Pi_{M}([p,v]))=\pi_{\hat{A}}([p,v])\quad\forall[p,v]\in P\times_{G}F$
@@
$\mathcal{L}_{\mathrm{GFM}}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ x\sim p_{t}\end{subarray}}\bigg\lVert\Big[v_{\theta}(x,t)-\alpha(t)\Pi_{M}(A_{\mu}(x,t)d^{\mu}(x,t)v_{s}(x,t))\Big]-u_{t}(x)\bigg\rVert_{g_{x}}^{2}$
@@ 4.1  Model
  • The mapping $\Pi_{M}:P\times_{G}\mathbb{R}^{N}\rightarrow M\times\mathbb{R}^{N}$ acts as the identity function on both $M$ and $\mathbb{R}^{N}$, with the $G$-component effectively projected out as a consequence of the chosen gauge.